OpenStudy (anonymous):

I'm having difficulty with orthogonal vectors, can anyone help me out? Find "a" so that the vectors v = i + aj and w = -4i - 2j are orthogonal.

OpenStudy (amistre64):

define orthogonal ....

OpenStudy (amistre64):

in slopes, we have orthogonal slopes (perpendicular) if one is the negative reciprocal of the other -4, -2 would be ortho to 2, -4 but 2,-4 isnt 1,a is it .. it is if we scale it :) divide by 2 1,-2 = 1,a

OpenStudy (amistre64):

otherwise dot and solve for a ...

OpenStudy (anonymous):

the orthogonal should be perpendicular.

OpenStudy (amistre64):

right and perpendicular slopes have a product of -1 a consequence to this is: if (a,b) and (x,y) are perpendular \[\frac ba \frac yx=-1\] \[\frac {by}{ax} =-1\] \[by =-ax\] \[by+ax=0\]

OpenStudy (amistre64):

the last line is just the dotproduct of the 2 vectors

OpenStudy (amistre64):

a, b . x, y ------- ax+by

OpenStudy (anonymous):

so in this case it would be: \[\frac{ a}{ 1}\frac{ 2 }{ 4} = -1\]

OpenStudy (anonymous):

4+2a = -1?

OpenStudy (amistre64):

v = i + aj and w = -4i - 2j a/1 and 2/4 a/2 = -1

OpenStudy (amistre64):

2a = -4 2a + 4 = 0

OpenStudy (anonymous):

\[\frac{ a }{ 2}=-1\] a = -2?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

the proper vector approach is to set the dotproduct to 0 1, a -4,-2 ----- -4-2a = 0 2a = -4 a = -4/2

OpenStudy (anonymous):

And that is it. wow, sometimes I way overthink this and make it too complicated.

OpenStudy (anonymous):

Thank you for your help! My finals week will be a little lighter thanks to you!

OpenStudy (amistre64):

good luck :)